OpenStudy (ihannah_banana7):

Evaluate the integral using the indicated trigonometric substitution. (Use C for the constant of integration.)

OpenStudy (ihannah_banana7):

\[\int\limits_{}^{}\frac{ x^3}{ \sqrt{x^2 +64} } dx\] \[x=8\tan \theta\]